7.(6 pts) Calculate the pH, [OH], and pOHl of a 0.0035 M HCl aqueous solution: 8. (6 pts) What are the [Hs01, (OH], and poH in a solution with a pH of 3.82? 9. (6 pts) What are the [H:0], [OH], and pH in a solution with a pOH of 11.75? pH = pOH= , [OH] = POH [OH] = 10. (6 pts) Calculate the [H30, [OH1, and pH a 0.040 M Ba(OH)2 solution. [OH-11- pH=
Answer #1

7) 0.0035M HCl

HCl -----> H+ + Cl-

[H+] = 0.0035M

Kw = 10-14 = [H+][OH-]

[OH-] = 10-14 / 0.0035 = 2.86 * 10-12M

pH = -log[H+] = -log(0.0035) = 2.45

pOH = -log[OH-] = -log(2.86 * 10-12) = 11.54

8) pH = 3.82

[H+] = 10-3.82 = 1.5 * 10-4 M

[OH-] = 10-14/ 1.5 * 10-4 = 6.67 * 10-11M

pOH = -log(6.67 * 10-11) = 10.17

9) pOH = 11.75

[OH-] = 10-11.75 = 1.77*10-12M

[H+] = 10-14 / 1.77*10-12 = 5.6*10-3 M

pH = -log(5.6*10-3) = 2.25

10) Ba(OH)2 ------> Ba2+ + 2OH-

[OH-] = 0.080M

[H+] = 10-14 / 0.080 = 1.25*10-13M

pH = -log(1.25*10-13) = 12.90